OpenStudy (anonymous):

Which set of numbers represents a Pythagorean triple? A. 6, 9, 12 B. 7, 10, 12 C. 16, 18, 25 D. 27, 36, 45

jhonyy9 (jhonyy9):

try like for example : A 6*6=36 added to 9*9 =81 so the last terms 6+1 need being equal 12*12=144 so the last term of 144 than you check it 7 not equal 4 so this mean A is fals B . 7*7=49 and 10*10=100 added get 149 and 12*12=144 so this mean fals too C. 16*16 result a number with end term of 6 and 18*18 result a number with end term of 4 so added 6+4 will result a number with end number 0 and you know that 25*25 result a number with end term of 5 so than this is false too D . hope this you can calcul it same like above